Sarah's recipe for 3 dozen raisin cookies but her cousin wants the recipe for 2 dozen. what is the recipe that sarah needs to send to marcus?
4 cups of sugar
1 cup of water
1 1/2 cup of shortening
2 cups of sugar
3 eggs
1/2 teaspoon of vanilla
4 cups of flour
1 teaspoon baking powder
1 teaspoon baking soda
2 teaspoons salt
1 1/4 teaspoons cinnamon
3/8 teaspoons nutmeg
3/8 teaspoons all spice
3/4 cup chopped pecans or walnuts

Answer by CharlesG2(834)

Sarah's recipe for 3 dozen raisin cookies but her cousin wants the recipe for 2 dozen. what is the recipe that sarah needs to send to marcus?
2 dozen = x * 3 dozen --> x=2/3 --> multiply all these ingredients by 2/3 and you will have recipe to send to Marcus
4 cups of sugar * 2/3 = 8/3 = 2 2/3 cups
1 cup of water * 2/3 = 2/3 cups
1 1/2 cup of shortening * 2/3 = 3/2 * 2/3 = 6/6 = 1 cup
2 cups of sugar * 2/3 = 4/3 = 1 1/3 cups
3 eggs * 2/3 = 2 eggs
1/2 teaspoon of vanilla * 2/3 = 2/6 = 1/3 teaspoon
4 cups of flour * 2/3 = 8/3 = 2 2/3 cups
1 teaspoon baking powder * 2/3 = 2/3 teaspoon
1 teaspoon baking soda * 2/3 = 2/3 teaspoon
2 teaspoons salt * 2/3 = 4/3 = 1 1/3 teaspoon
1 1/4 teaspoons cinnamon * 2/3 = 5/4 * 2/3 = 20/12 = 10/6 = 5/3 = 1 2/3 teaspoon
3/8 teaspoons nutmeg * 2/3 = 6/24 = 1/4 teaspoon
3/8 teaspoons all spice * 2/3 = 1/4 teaspoon
3/4 cup chopped pecans or walnuts * 2/3 = 6/12 = 1/2 cup